OpenStudy (anonymous):

graph the equation by plotting three points -3y=x-6

OpenStudy (jdoe0001):

same way as the one at http://openstudy.com/study#/updates/51fd8387e4b0cc46c14aa80c

OpenStudy (anonymous):

I don't what three points to use for x

OpenStudy (phi):

can you divide both sides by -3 (to get y by itself) ?

OpenStudy (anonymous):

i have to find three points for x

OpenStudy (phi):

yes, and to make it easier to do, we want to simplify your equation we want y = mx + b you have -3y = x-6 to get y by itself, divide both sides and all terms by -3 can you do that ?

OpenStudy (anonymous):

1y=x+2

OpenStudy (phi):

almost. -3y / -3 = x/-3 + -6/-3 you get \[ 1y= \frac{1}{-3} x + 2\] or just \[ y= \frac{1}{-3} x + 2\]

OpenStudy (phi):

now if x = 0 , what is y ? use the equation

OpenStudy (anonymous):

0,2

OpenStudy (phi):

so that is your first point the slope is 1/-3 or -1/3 (both work) the slope is change in y divided by change in x if you change x by +3 , y changes by -1 can you find the next point ?

OpenStudy (anonymous):

3,1

OpenStudy (phi):

yes. now the next point ?

OpenStudy (anonymous):

6,0

OpenStudy (phi):

yes. notice that if you use a slope of 1/-3 the change in y divided by change in x change in y is +1 and the change in x is -3 so if you started at (0,2) and added -3 on the x, you would add +1 to the y. you get (-3,3) as a point on the line

OpenStudy (anonymous):

yes its so much easier that way

OpenStudy (anonymous):

thank you